A flask contains a mixture of Nitrous Oxide(N2O) and Oxygen(O2) with partial pressures of 702 torr and 749 torr.
What is the mole ratio of the number of moles of N2O to the number of moles of O2 in this flask?
1
Expert's answer
2019-10-02T08:03:36-0400

Due to Dalton law and Ideal gas law

Gas partial pressure is equal to mole ratio.

Then:

P = 702 + 749 = 1451 torr

x(O2) = 749/1451 = 0.516

x(N2O) = 1 - 0.516 = 0.484

N2O/O2 = 0.484/0.516 = 0.938